OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 15 meters and legs that measure 9 meters and 12 meters. What is the surface area of this prism if the height of the prism is 11 meters? 450 m2 486 m2 504 m2 612 m2

OpenStudy (skullpatrol):



OpenStudy (anonymous):

The question didnt give me a drawing

OpenStudy (skullpatrol):

You are expected to make one yourself.

OpenStudy (anonymous):

no its a mutiple choice answer

OpenStudy (skullpatrol):

That doesn't matter.

OpenStudy (anonymous):

It never says that i have to draw a figure

OpenStudy (skullpatrol):

Can you visualize a prism?

OpenStudy (skullpatrol):

Do you want to visualize this prism?

OpenStudy (anonymous):

i dont unnderstand this

OpenStudy (skullpatrol):

Do you understand what surface area is?

OpenStudy (skullpatrol):

What is the surface area of this prism if the height of the prism is 11 meters?

OpenStudy (skullpatrol):

Surface area is the same as regular area just extended to a 3 dimensional object.

OpenStudy (skullpatrol):

In this case you have 5 surface areas to calculate and then add up to give the total surface area.

OpenStudy (skullpatrol):

Are you familiar with the formula to calculate the area of a rectangle?

Directrix (directrix):

@tj_1502 The surface area of the prism can be found by the following formula: SA = p*h + 2*B where p is the perimeter of the base and B is the area of the base. SA = (15+9+12)*11 + 2*(1/2)*9*12 SA = ?

OpenStudy (anonymous):

Directrix can you help me please

OpenStudy (skullpatrol):

So you don't want my help?

OpenStudy (anonymous):

i just dont get anything your trying to tell me

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):

Study this sketch, and try to think about the "surface area" of it.

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

Do you know the definition of area?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

What is it?

Directrix (directrix):

@tj_1502 SA = (15+9+12)*11 + 2*(1/2)*9*12 SA = 36*11 + 9*12 SA = ?